Formulate a system of equations for the situation below and solve.
The Johnson Farm has 400 acres of land allotted for cultivating corn and wheat. The cost of cultivating corn and wheat (including seeds and labor) is $48 and $36 per acre, respectively. Jacob Johnson has $18,600 available for cultivating these crops. If he wishes to use all the allotted land and his entire budget for cultivating these two crops, how many acres of each crop should he plant?

Answer by richwmiller(17219)

x+y=400,
48*x+36*y=18600
x=400-y
48*(400-y)+36*y=18600
19200-48y+36*y=18600
-12*y=-600
y=50 at $36
x=400-y
x=350 at $8
check
48*350+036*50=18600
16800+1800=18600
18600=18600
ok